# Tree of Attributes Prompt Learning for Vision-Language Models

### v1

@@ Abstract
Prompt learning has proven effective in adapting vision language models for downstream tasks. However, existing methods usually append learnable prompt tokens solely with the category names to obtain textual features, which fails to fully leverage the rich context indicated in the category name. To address this issue, we propose the Tree of Attributes Prompt learning (TAP), which first instructs LLMs to generate a tree of attributes with a “concept - attribute - description” structure for each category, and then learn the hierarchy with vision and text prompt tokens. Unlike existing methods that merely augment category names with a set of unstructured descriptions, our approach essentially distills structured knowledge graphs associated with class names from LLMs. Furthermore, our approach introduces text and vision prompts designed to explicitly learn the corresponding visual attributes, effectively serving as domain experts. Additionally, the general and diverse descriptions generated based on the class names may be wrong or absent in the specific given images. To address this misalignment, we further introduce a vision-conditional pooling module to extract instance-specific text features. Extensive experimental results demonstrate that our approach outperforms state-of-the-art methods on the zero-shot base-to-novel generalization, cross-dataset transfer, as well as few-shot classification across 11 diverse datasets.
@@ 1 Introduction
In parallel, another stream of research has been exploring the utilization of large language models (LLMs) to generate more elaborate and descriptive text prompts for enhancing zero-shot learning capabilities (Menon & Vondrick, 2023; Pratt et al., 2023; Roth et al., 2023; Kim et al., 2023; Parkhi et al., 2012; Yan et al., 2023; Yang et al., 2023; Roy & Etemad, 2024; Zheng et al., 2023; Tian et al., 2023). These LLM-generated descriptions offer a wealth of detail and context, potentially enriching the model’s interpretative capabilities. However, current methodologies in integrating these descriptions often do not exploit the full potential of this richness. As shown in Fig. Fig. 1 (b), most of these approaches lack a structured framework to organize and utilize these descriptions effectively, leading to a scattergun approach where not all generated descriptions are contextually relevant or optimally aligned with the visual content. In addition, as noted in (Roth et al., 2023), descriptions generated by such paradigms are usually diverse, which covers most possibilities of the class, but include descriptions that are either likely not co-occurring, e.g. “steamed” and “fried”, or absent in the input image, e.g. “long tail” for a cat shot from the front, necessitating the need for a selective pooling mechanism for clearer image-text alignments.
Figure 1: Illustration of the methods for CLIP text prompts formation. (a) Manually created prompt with the single “a photo of a {class}” template; (b) A unstructured set of detailed descriptions generated by LLMs; (c) The proposed Tree of Attribute that organizes the descriptions in a “concept - attribute - descriptions” structure, essentially distilling knowledge graphs from LLMs; (d) An example Tree of Attribute for “dumplings”.
@@ 2 Related Work
Prompt Learning for Vision-Language Models. Prompt learning bridges linguistic understanding and visual perception by guiding VLMs with text prompts, a concept originated in NLP (Lester et al., 2021; Li & Liang, 2021; Liu et al., 2021) and adapted to vision-only (Jia et al., 2022; Wang et al., 2022a; b; Zhang et al., 2022) and multimodal contexts(Zhou et al., 2022b; a; Khattak et al., 2023a; b; Shi & Yang, 2023; Lee et al., 2023; Tian et al., 2023; Rasheed et al., 2023; Roy & Etemad, 2024; Zheng et al., 2023; Zhu et al., 2023; Bulat & Tzimiropoulos, 2023; Lu et al., 2022). In the textual domain, CoOp (Zhou et al., 2022b) optimizes learnable continuous prompts in CLIP’s language branch for few-shot image recognition, while CoCoOp (Zhou et al., 2022a) addresses CoOp’s overfitting issues by conditioning prompts on visual features. In the visual domain, Visual Prompt Tuning (VPT) (Bahng et al., 2022) and Dual-modality Prompt Tuning (DPT) (Xing et al., 2023) enhance CLIP’s vision encoder by learning visual prompts in pixel space and dynamically generating prompts through cross-attention, respectively. TransHP (Wang et al., 2023b) leverages category hierarchy for prompt learning to improve classification performance. LoGoPrompt (Shi & Yang, 2023) enhances classification by incorporating synthetic images with class name text as auxiliary visual prompts. MaPLe (Khattak et al., 2023a) explores multimodal prompt learning, jointly optimizing prompts in both vision and language branches. Other recent works have focused on regularizing prompt learning to leverage the knowledge from base VLMs effectively, demonstrating enhanced generalization in varied downstream visual tasks (Khattak et al., 2023b; Bulat & Tzimiropoulos, 2023; Roy & Etemad, 2024). PromptSRC, for instance, introduced a self-regulating method that restricts both the vision and text prompt, demonstrating improved generalization. Distinct from these approaches, PLOT (Chen et al., 2023) and ALIGN (Wang et al., 2023a) leverage Optimal Transport to align multiple prompts with local visual features, either from the multi-head self-attention layer or at a token level. Our work diverges from these methods by introducing a hierarchical "Tree of Attribute" framework derived from LLMs to structure textual descriptions and guide the learning of specialized "domain expert" tokens for attribute-level understanding.
Enhancing model’s understanding using visual attributes. There’s a growing emphasis on enhancing model’s understanding of the class names by integrating visual attributes into the model. A common to do that is to use detailed visual descriptions from LLMs, moving beyond generic prompts (Zhou et al., 2022b; a). These descriptions, like the “fur pattern” or “tail shape” of a cat, provide fine-grained and distinctive characteristics. The use of LLMs like GPT-3 (Brown et al., 2020), allows for efficient generation of a broad spectrum of class-specific descriptions, offering an advantage over manually crafted templates. While this approach has been extensively researched in zero-shot contexts (Kim et al., 2023; Menon & Vondrick, 2023; Parkhi et al., 2012; Roth et al., 2023; Yan et al., 2023; Yang et al., 2023; Fabian et al., 2023; Pratt et al., 2023; Novack et al., 2023), its application in conjunction with prompt learning for few-shot tasks remains relatively unexplored(Mao et al., 2023; Lee et al., 2023; Tian et al., 2023; Zheng et al., 2023; Zhang et al., 2024; Liu et al., 2024). Additionally, another line of research integrates the visual attributes through data augmentations. For example, AAPL (Kim et al., 2024) uses adversarial token embedding to mitigate bias and enhance attribute-specific learning through augmented image features. However, we note the following limitations of previous methodologies. First, in the use of LLMs, existing methods have largely utilized unstructured descriptions, which lacks an organized framework for effective utilization. Additionally, while data augmentation can improve learning of invariant visual attributes, it is constrained by the limitations of image data augmentation, which cannot cover all possible variations. Our approach diverges by structuring the visual descriptions into a “Tree of Attribute”, coupled with learnable visual prompts as domain experts. Additionally, LLM-generated descriptions often cover a wide range of potential class descriptions, of which not all may be pertinent to a given image, pointing to the need for a selective pooling mechanism to ensure optimal image-text alignment. We further introduce a vision-conditional pooling layer for refined image-text alignment. This structured approach not only enhances the interpretability of the model’s learning process but also significantly improves alignment accuracy between image content and descriptive text.
@@ 3.6 Training and Inference
Table 1: Comparison with state-of-the-art methods in base-to-novel generalization. HM: harmonic mean (Xian et al., 2017).
@@ 4.2 Cross-dataset Transfer
Table 2: Comparison with state-of-the-art methods in cross-dataset transfer evaluation.
@@ 4.4 Ablation Study
Effects of Tree of Attribute. A core inquiry is whether structuring descriptions into a Tree of Attribute (ToA) offers advantages over an unstructured aggregation of LLM-generated descriptions. To evaluate, we revert to aligning a mixed, unstructured set of descriptions with the CLS token - a common practice in prior studies (Mao et al., 2023; Lee et al., 2023; Tian et al., 2023; Zheng et al., 2023), while keeping the same number of visual prompt tokens. According to Table 4, substituting the ToA with an unstructured set results in significant performance decreases of 1.86%, 2.31%, and 2.11% across the average base, novel, and their harmonic mean performances, respectively. This stark contrast underscores the ToA’s critical role in enhancing model efficacy.
@@ 4.5 Visualization
Table 8: Design choice of the pooling layer.

### v2

@@ Abstract
Prompt learning has proven effective in adapting vision language models for downstream tasks. However, existing methods usually append learnable prompt tokens solely with the category names to obtain textual features, which fails to fully leverage the rich context indicated in the category name. To address this issue, we propose the Tree of Attributes Prompt learning (TAP), which first instructs LLMs to generate a tree of attributes with a “concept - attribute - description” structure for each category, and then learn the hierarchy with vision and text prompt tokens. Unlike existing methods that merely augment category names with a set of unstructured descriptions, our approach essentially distills structured knowledge graphs associated with class names from LLMs. Furthermore, our approach introduces text and vision prompts designed to explicitly learn the corresponding visual attributes, effectively serving as domain experts. Additionally, the general and diverse descriptions generated based on the class names may be wrong or absent in the specific given images. To address this misalignment, we further introduce a vision-conditional pooling module to extract instance-specific text features. Extensive experimental results demonstrate that our approach outperforms state-of-the-art methods on the zero-shot base-to-novel generalization, cross-dataset transfer, as well as few-shot classification across 11 diverse datasets. Code is available at https://github.com/HHenryD/TAP.
@@ 1 Introduction
In parallel, another stream of research has been exploring the utilization of large language models (LLMs) to generate more elaborate and descriptive text prompts for enhancing zero-shot learning capabilities (Menon & Vondrick, 2023; Pratt et al., 2023; Roth et al., 2023; Li et al., 2024; Kim et al., 2023; Parkhi et al., 2012; Yan et al., 2023; Yang et al., 2023; Roy & Etemad, 2024; Zheng et al., 2023; Tian et al., 2023). These LLM-generated descriptions offer a wealth of detail and context, potentially enriching the model’s interpretative capabilities. However, current methodologies in integrating these descriptions often do not exploit the full potential of this richness. As shown in Fig. 1 (b), most of these approaches lack a structured framework to organize and utilize these descriptions effectively, leading to a scattergun approach where not all generated descriptions are contextually relevant or optimally aligned with the visual content. In addition, as noted in (Roth et al., 2023), descriptions generated by such paradigms are usually diverse, which covers most possibilities of the class, but include descriptions that are either likely not co-occurring, e.g. “steamed” and “fried”, or absent in the input image, e.g. “long tail” for a cat shot from the front, necessitating the need for a selective pooling mechanism for clearer image-text alignments.
Figure 1: Illustration of the methods for CLIP text prompts formation. (a) Manually created prompt with the single “a photo of a {class}” template; (b) A unstructured set of detailed descriptions generated by LLMs; (c) The proposed Tree of Attribute distills a knowledge graph from LLMs, organizing the knowledge in “concept - attribute - descriptions” structure; (d) An example Tree of Attribute for class “dumplings”, where each color represents a visual attribute.
@@ 2 Related Work
Prompt Learning for Vision-Language Models. Prompt learning bridges linguistic understanding and visual perception, originating in NLP (Lester et al., 2021; Li & Liang, 2021; Liu et al., 2021) and later adapted to vision-only (Jia et al., 2022; Wang et al., 2022a; b; Zhang et al., 2022) and multimodal settings (Zhou et al., 2022b; a; Khattak et al., 2023a; b; Shi & Yang, 2023; Lee et al., 2023; Tian et al., 2023; Rasheed et al., 2023; Roy & Etemad, 2024; Zheng et al., 2023; Zhu et al., 2023; Bulat & Tzimiropoulos, 2023; Lu et al., 2022). CoOp (Zhou et al., 2022b) introduced learnable continuous prompts for few-shot image recognition but struggled to generalize to unseen classes, highlighting the challenge of base-to-novel generalization (Zhou et al., 2022a; Guo et al., 2024b; Hernandez-Hernandez et al., 2024; Guo et al., 2024a). CoCoOp (Zhou et al., 2022a) addressed this by conditioning prompts on visual features. Visual and multimodal prompt tuning methods, such as VPT (Bahng et al., 2022), DPT (Xing et al., 2023), and MaPLe (Khattak et al., 2023a), optimize prompts in pixel or text space to enhance feature alignment. Other works (Khattak et al., 2023b; Bulat & Tzimiropoulos, 2023; Li et al., 2022; Roy & Etemad, 2024) focus on regularization to improve generalization. PromptSRC introduced self-regulating prompts to better retain base knowledge, while methods like PLOT (Chen et al., 2023) and ALIGN (Wang et al., 2023) apply Optimal Transport to align prompts with local features. Our work differs by introducing a hierarchical "Tree of Attribute" framework to structure textual descriptions and guide specialized "domain expert" tokens for attribute-level understanding.
Enhancing model’s understanding using visual attributes. There’s a growing emphasis on the use of detailed visual descriptions for various visual understanding tasks, including more fine-grained captioning (Hsieh et al., 2024), identifying subordinate-level categories (Liu et al., 2024a), and language-guided visual classification (Menon & Vondrick, 2023). While manual creation is impractical given the large number of image classes, existing research relies either on data augmentation (Kim et al., 2024) or generation by LLMs such as GPT-3 (Brown et al., 2020), which offers an efficient generation of a broad spectrum of class-specific descriptions. These descriptions, like “fur pattern” or “tail shape” of a cat, provide fine-grained and distinctive characteristics. In an essence, such approaches can be viewed as knowledge distillation from LLMs trained on trained on vast and diverse textual corpora. However, existing studies often lack a structured methodology for distillation (Kim et al., 2023; Menon & Vondrick, 2023; Parkhi et al., 2012; Roth et al., 2023; Yan et al., 2023; Yang et al., 2023; Fabian et al., 2023; Pratt et al., 2023; Novack et al., 2023; Mao et al., 2023; Tian et al., 2023; Zheng et al., 2023; Zhang et al., 2024; Liu et al., 2024b) or fail to effectively exploit the inherent hierarchy within the knowledge (Maniparambil et al., 2023; Wang et al., 2024; Hsieh et al., 2024; Liu et al., 2024a). Our approach (TAP ) addresses these limitations by introducing a novel method to distill a knowledge graph from LLMs in a top-down manner, transitioning from class names (concepts) to visual attributes (e.g., color, shape) and further to detailed descriptions of each attribute, forming a structured Tree of Attributes (ToA). To fully leverage the ToA, we propose a bottom-up integration pipeline. We introduce vision-conditional pooling (VCP) layers to aggregate descriptions into attribute-level features, effectively mitigating potential noise in the generated descriptions. The alignment between attributes and introduced visual expert tokens is then refined through this hierarchical structure. This integration enables the model to exploit structured relationships within the ToA, enhancing both the granularity and interpretability of vision-text alignment.
@@ 3.6 Training and Inference
Table 1: Comparison with state-of-the-art methods in base-to-novel generalization. The model is trained on the base class, and evaluated on the unseen novel classes in zero-shot. TAP demonstrates strong generalization performance. HM: harmonic mean (Xian et al., 2017).
@@ 4.2 Cross-dataset Transfer
Table 2: Comparison with state-of-the-art methods in cross-dataset transfer evaluation. The model is trained on the source dataset and evaluated on the target datasets in zero-shot.
@@ 4.4 Ablation Study
Effects of Tree of Attribute. A core inquiry is whether structuring descriptions into a Tree of Attribute (ToA) offers advantages over an unstructured aggregation of LLM-generated descriptions. To evaluate, we revert to aligning a mixed, unstructured set of descriptions with the CLS token - a common practice in prior studies (Mao et al., 2023; Maniparambil et al., 2023; Liu et al., 2024b; Wang et al., 2024; Tian et al., 2023; Zheng et al., 2023), while keeping the same number of visual prompt tokens. According to Table 4, substituting the ToA with an unstructured set results in significant performance decreases of 1.86%, 2.31%, and 2.11% across the average base, novel, and their harmonic mean performances, respectively. This stark contrast underscores the ToA’s critical role in enhancing model efficacy.
@@ 4.5 Visualization
Table 8: Ablation on design choice of the VCP layer. Our cross-attention based pooling mechanism demonstrates the best performance among other variants.
@@ 5 Conclusion
Acknowledgment
This work was supported in part by Microsoft’s AI for Good Research Lab, the Harvard Data Science Initiative, and NIH Grant R01HD104969.
@@ A.4 Prompts for Tree-of-Attribute generation
A.5 Attribute sets
The attribute sets generated by LLMs are shown in Table 11 - 12.
Table 11: Attribute sets generated by LLMs for the 11 datasets.
Dataset Attributes
ImageNet Orientation
Shape
Pattern
Texture
Pose
Context
Caltech101 Dominant Feature
Shape
Texture
Color
Size
StanfordCars Body Type
Wheel Design
Grille Style
Headlight Shape
Rear Taillight Design
Roof Style
Flowers102 Color
Petal
Center structure
Stem characteristics
Food101 Color
Shape
Texture
Ingredients
Presentation Style
Table 12: Attribute sets generated by LLMs for the 11 datasets. Cont.
Dataset Attributes
FGVCAircraft Wing Configuration
Winglet Presence
Engine Configuration
Number of Engines
Fuselage Length
Fuselage shape
Wingspan
SUN397 Indoor/Outdoor
Color
Dominant elements
Environment
Architectural style
Patterns
DTD Texture
Pattern
Repetition
Contrast
EuroSAT Contrast
Texture
Orientation
Edge
Size
Color
Symmetry
UCF101 Action Pose
Number of People
Background Setting
Objects Present